OpenStudy (anonymous):

Find 3 consecutive odd numbers where the product of the smaller two numbers is 22 less than the square of the largest number.

OpenStudy (anonymous):

@sid0830

OpenStudy (johnweldon1993):

Let the first number = 2x + 1 The seconds is 2x + 3 The third is 2x + 5 The product of the 2 smaller numbers is 22 less than the square of the third...so we have \[\large (2x + 1)(2x + 3) = (2x + 5)^2- 22\]

OpenStudy (anonymous):

x(x+2)=(x+4)^2-22

OpenStudy (anonymous):

\[(x)(x+2)=(x+4)^2−22\] x=1 ..answer :)

OpenStudy (anonymous):

n=1